NEXT
Make New Entry,
Make Followup Entry
User name riordan
Log entry time 18:04:31 on March 08, 2009
Entry number 265340
keyword=Summary for March 7 Swing - March 8 Day
Summary for Mar 7 swing - Mar 8 Day
Finished longitudinal running and 4I. Started transverse and 4G.
Day shift characterized by ioc crashes and Left Q1 and Dipole levels
tripping the magnets.
1L04 vacuum level continues to rise throughout the day.
Average transverse target polarization: 58.2%
Beam in Use
Shift |
ABU |
BANU |
BNA |
2009/03/07 Swing | 3.95 |
1.37 | 2.68 |
2009/03/08 Owl | 4.49 |
0.75 | 1.76 |
2009/03/08 Day | 4.31 |
1.25 | 2.37 |
TOTAL | 12.75 |
3.37 | 6.81 |
Main events/issues:
- Mar 7
- Finished longitudinal running. Moved to transverse.
- Noticed vacuum in 1L04 was rising
- Mar 8
- 7:40 iochla and a few other iocs go down and we cannot be
supplied beam
- 9:30 iocs back up, left Q1 trips
- Periodic drops in cryo levels in left Q1 and left dipole occur
due to the fact that one ioc was not set to the proper state
- 17:15 ioc restored, levels look stable
- 1L04 vacuum continues to degrade
Plans for the next few days:
- Half wave plate change Sunday night
- Continue 4-pass transverse production
- Monday day shift will be positive polarity running if vacuum behaves
- Tuesday beam studies will involve opening beamline vacuum to
install material around the compton electron detector.
A copy of this log entry has been emailed to: brads, jpchen, jiang, meziani, choi, riordan